# Cross-composition Feature Disentanglement for Compositional Zero-shot Learning

### v1

@@ Abstract
Disentanglement of visual features of primitives (i.e., attributes and objects) has shown exceptional results in Compositional Zero-shot Learning (CZSL). However, due to the feature divergence of an attribute (resp. object) when combined with different objects (resp. attributes), it is challenging to learn disentangled primitive features that are general across different compositions. To this end, we propose the solution of cross-composition feature disentanglement, which takes multiple primitive-sharing compositions as inputs and constrains the disentangled primitive features to be general across these compositions. More specifically, we leverage a compositional graph to define the overall primitive-sharing relationships between compositions, and build a task-specific architecture upon the recently successful large pre-trained vision-language model (VLM) CLIP, with dual cross-composition disentangling adapters (called L-Adapter and V-Adapter) inserted into CLIP’s frozen text and image encoders, respectively. Evaluation on three popular CZSL benchmarks shows that our proposed solution significantly improves the performance of CZSL, and its components have been verified by solid ablation studies. Our code and data are available at: https://anonymous.4open.science/r/DCDA-0BF7 .
Introduction
Compositional Zero-shot Learning (CZSL), which learns visual primitives from seen compositions of attributes and objects and combines them to recognize novel unseen compositions, has attracted extensive research interest in the past few years (Misra, Gupta, and Hebert 2017). For example, when presented with images of compositions like red tomato and green apple, the model could instantly recognize the new composition green tomato by combining the attribute green with the object tomato, even though it has never seen images of green tomato. This ability not only enables a deep neural network to tell plenty of new concepts with zero examples (Chen et al. 2023), but also facilitates the study of vision-language understanding (Chen et al. 2024) as the image features need to be accurately decomposed and recomposed with the guidance of textual attribute-object labels.
One solution for CZSL is to build a shared embedding space where the image features and the composition embeddings can be compared to perform classification (Wei et al. 2019; Naeem et al. 2021; Mancini et al. 2022). Recently, its performance has been greatly improved by integrating large pre-trained vision-language models like CLIP (Radford et al. 2021) which jointly pre-trains an image encoder and a text encoder with visual-semantic alignment (Nayak, Yu, and Bach 2023). However, attributes and objects often highly depend on each other in images and their encoded visual features are too entangled to decompose. Red tomato is a typical example whose image pixels for red and tomato totally overlap. In view of this, recent CLIP-based methods either design separated soft prompts to learn disentangled textual features for attributes and objects and coax the model into identifying the single primitive in an image (Lu et al. 2023; Wang et al. 2023a), or insert trainable and parameter-independent adapters into the frozen image encoder to learn disentangled visual primitive features directly (Zheng, Zhu, and Nevatia 2024; Huang et al. 2024; Li et al. 2024). These disentangled representations enable finer-grained visual-semantic alignment at the primitive level.
However, the above methods treat the disentanglement of each composition independently, ignoring the diversity of attributes and objects across different compositions, i.e., an attribute (resp. object) tends to look greatly different when paired with different objects (resp. attributes); for example, in the compositions of red apple, red tomato and red wine, the attribute red is diverse e.g., the red tone ranges from shallow to deep. Such diversity may make the learned primitive features especially the visual ones less discriminative, thus limiting their generalization to their unseen compositions. In the left of Figure 1, we randomly sample some testing images whose annotated attributes have high diversity from the MIT-States dataset (Isola, Lim, and Adelson 2015), and visualize their disentangled attribute features of images learned by the previous state-of-the-art (SOTA) method CAILA (Zheng, Zhu, and Nevatia 2024). The disentangled features of images that share the same attribute mostly scatter in the vector space with limited discrimination. For example, the features of broken (purple circles) are fully mixed with the features of cooked, coiled, mossy and so on.
To obtain the disentangled primitive features that are more discriminative and general across different compositions (as in the right of Figure 1), we argue that introducing other compositions having the same attribute or object as the target composition during feature disentanglement is beneficial. Inspired by (Naeem et al. 2021), we propose to utilize a compositional graph to define the overall primitive-sharing relationships between compositions, where all the attributes, objects, and compositions act as graph nodes and their compositional relationships are modeled by edges, as shown in the bottom of Figure 2(a). Based on this graph, we thereby propose Dual Cross-composition Feature Decomposing Adapters (DCDA) to augment the CLIP encoders on both language and vision sides so as to learn disentangled textual and visual representations of each primitive that are cross-composition generalizable.
Since the textual primitive features are less entangled than the visual primitive features due to the separated words in the label, DCDA has different implementations for the language and vision sides. Specifically, the adapter on the language side, called L-Adapter, applies a graph neural network (GNN) to the above compositional graph, to propagate the encoded textual features of neighboring compositions to an attribute or object node for its cross-composition generalizable primitive representation. Inspired by (Hao, Han, and Wong 2023), the adapter on the vision side, called V-Adapter, is first built upon the cross-attention between images of two primitive-sharing compositions (e.g., red tomato and red wine) to highlight the image features highly relevant to each other, i.e., the cross-composition generalizable features of their common primitive (e.g., red). Then, we propose a novel primitive relevance-guided sampling strategy to effectively introduce more primitive-sharing compositions for more general primitive features, where two primitive graphs are extracted from the compositional graph to describe the attribute and object relevance. When integrating L-Adapters and V-Adapters into multiple layers of CLIP’s text and image encoders, we retain the original parameters of CLIP to avoid overfitting, but inject the task-specific knowledge. Our contributions can be summarized below:
  • DCDA is among the first to investigate cross-composition feature disentanglement in CLIP-based CZSL, with the diversity of primitives concerned.
  • A compositional graph and two derived primitive graphs are utilized to define exhaustive primitive-sharing compositions for less entangled text labels and more entangled image samples, respectively, with two kinds of effective adapters accordingly designed.
  • DCDA achieves new SOTA results on MIT-States and UT-Zappos benchmarks in both closed world and open world, and has very competitive performance on C-GQA. Compared with CAILA, our method disentangles more clustered, more distinctive, and more generalizable attribute embeddings, as shown in the right of Figure 1. More details and comparisons on the disentangled object embeddings are attached in Appendix C.
@@ Introduction
Figure 1: $t$-SNE visualizations of disentangled attribute representations of images in the test set of MIT-States, learned by CAILA (Zheng, Zhu, and Nevatia 2024) and our DCDA. Solid and hollow circles represent images of seen and unseen compositions, respectively. Best viewed in color.
Related Work
@@
Conventional CZSL methods are roughly divided into two groups. One is classifier-based which first trains two separate classifiers to predict an input image’s attribute and object labels, respectively, and then combines them to predict the compositional labels (Misra, Gupta, and Hebert 2017; Nagarajan and Grauman 2018). The subsequent works further propose to enhance the dependence of the attribute and object in a composition (Li et al. 2020, 2022; Wang et al. 2023b). The other group is embedding-based which jointly represents attributes and objects to capture the dependence, and then aligns them with the images in a shared embedding space (Wei et al. 2019; Karthik, Mancini, and Akata 2022). In particular, (Naeem et al. 2021) learn the joint representation through graph convolutional networks. There are also some works concerning the disentanglement of attribute and object features in the visual space (Saini, Pham, and Shrivastava 2022; Hao, Han, and Wong 2023; Kim et al. 2023). However, all these methods have to learn the alignment between image features and text embeddings from scratch and are prone to overfit to the seen compositions.
@@
CLIP-based CZSL. After pre-training using 400 million image-text pairs, CLIP can be applied to any visual classification task without fine-tuning by setting prompts like “a photo of [class]”, where “[class]” is filled with the name of the class to be recognized. (Nayak, Yu, and Bach 2023) then had the first attempt to design prompt “a photo of [attribute] [object]” for CZSL, where “[attribute] [object]” are tunable tokens to teach CLIP how to compose attributes and objects.
@@
To stress the roles of individual primitives, HPL (Wang et al. 2023a) additionally sets an attribute and an object prompt with only “[attribute]” or “[object]” tunable; DFSP (Lu et al. 2023) makes the whole prompt trainable and fuses the decomposed text features with the encoded (entangled) image features through a cross-modal fusion module. Different from these works focusing on optimizing the prompts, CAILA (Zheng, Zhu, and Nevatia 2024) proposes to insert trainable adapters inside the frozen transformer layers to decompose and recompose the attribute and object features. With disentangled primitive features, Troika (Huang et al. 2024) establishes three prediction branches and pulls a static class prompt to its dynamic images via a cross-modal traction module. (Li et al. 2024) investigate the relative specificity of attributes when paired with different objects. In contrast to these methods, our method DCDA is more generalizable, with cross-composition knowledge injected and dual adapters inserted in CLIP’s image and text encoders.
@@
Methodology
CZSL Task Formulation. Let $\mathcal{D}_{tr}=\{(x,c)|x\in\mathcal{X}_{s},c\in\mathcal{C}_{s}\}$ be the training set, where $\mathcal{X}_{s}$ contains the training images and $\mathcal{C}_{s}$ is a set of seen compositional labels that are available during training. Each label is a tuple $c=(a,o)$ of an attribute class $a\in\mathcal{A}$ and an object class $o\in\mathcal{O}$. After training, the CZSL model can predict images of a set of new compositions $\mathcal{C}_{u}$ that are unseen during training, with $\mathcal{C}_{u}\cap\mathcal{C}_{s}=\emptyset$. Following previous works, we study generalized CZSL (Purushwalkam et al. 2019), where images of seen and unseen compositions are both tested and the candidate label space includes both seen and unseen labels. The test set is thus denoted as $\mathcal{D}_{te}=\{(x,c)|x\in\mathcal{X}_{te},c\in\mathcal{C}_{te}\}$, where $\mathcal{X}_{te}=\mathcal{X}_{u}\cup\mathcal{X}^{\prime}_{s}$ with $\mathcal{X}^{\prime}_{s}\cap\mathcal{X}_{s}=\emptyset$, and $\mathcal{C}_{te}=\mathcal{C}_{u}\cup\mathcal{C}_{s}$. Notably, $\mathcal{C}_{s}$ and $\mathcal{C}_{u}$ share the same attribute set $\mathcal{A}$ and object set $\mathcal{O}$, CZSL assumes that each $a$ and $o$ has been trained before testing and only the composition $(a,o)\in\mathcal{C}_{u}$ is novel.
@@
We first define the compositional graph. It consists of $N=|\mathcal{A}|+|\mathcal{O}|+|\mathcal{C}^{\prime}|$ nodes, including all the attributes, all the objects, and the compositions in the current computation (i.e., $\mathcal{C}^{\prime}=\mathcal{C}_{s}$ for training and $\mathcal{C}^{\prime}=\mathcal{C}_{te}$ for testing). Given these nodes, for each $c=(a,o)$, we connect $(a,o)$, $(a,c)$ and $(o,c)$ to form a triangle in the graph, as shown in Figure 2(a). For simplicity and efficiency, we keep all graph edges unweighted and undirected as in (Naeem et al. 2021) and obtain a symmetric adjacency matrix $A\in\mathbb{R}^{N\times N}$ to store the graph structure, $A_{ij}=1$ if there is a connection between node $i$ and $j$ otherwise $A_{ij}=0$.
@@
Next, we exploit multiple GNN layers to propagate features among graph nodes following the graph structure defined in $A$. Since many typical GNN models such as GCN (Kipf and Welling 2017), GAT (Velickovic et al. 2018) can be used here, we illustrate this procedure with two functions commonly defined in GNNs. One is agg for aggregating features of one-hop neighbors for each node, and the other is con for fusing the node feature and the neighborhood feature to make an update. Formally, as Figure 2(a)’s top shows, for each $c=(a,o)$, three identical agg functions are parallelly applied to aggregate neighborhood features for nodes $a,o,c$ at the $k$-th GNN layer, $k\in\{0,...,K-1\}$, as:
$\bm{a}^{(k)}_{\mathcal{N}_{a}}=\textsc{agg}^{(k)}(\{\bm{c}^{(k)}_{i}|c_{i}\in\mathcal{N}_{a}^{c}\},\{\bm{o}^{(k)}_{i}|o_{i}\in\mathcal{N}_{a}^{o}\})$ (1)
$\bm{o}^{(k)}_{\mathcal{N}_{o}}=\textsc{agg}^{(k)}(\{\bm{c}^{(k)}_{j}|c_{j}\in\mathcal{N}_{o}^{c}\},\{\bm{a}^{(k)}_{j}|a_{j}\in\mathcal{N}_{o}^{a}\})$ (2)
$\bm{c}^{(k)}_{c=(a,o)}=\textsc{agg}^{(k)}(\bm{a}^{(k)},\bm{o}^{(k)})$ (3)
where $\mathcal{N}_{a}^{c}$ (resp. $\mathcal{N}_{o}^{c}$) denotes the composition neighbor set of $a$ (resp. $o$) on the graph, and $\mathcal{N}_{a}^{o}$ (resp. $\mathcal{N}_{o}^{a}$) includes the objects (resp. attributes) that compose the compositions in $\mathcal{N}_{a}^{c}$ (resp. $\mathcal{N}_{o}^{c}$) together with $a$ (resp. $o$). Considering the example in Figure 2(a) where $c=(a,o)$ is red tomato, $\mathcal{N}_{a}^{c}$ includes compositions like red apple and $\mathcal{N}_{a}^{o}$ includes objects like apple. While the neighbor set of each $c$ only contains its primitives $a$ and $o$. $\bm{a}^{(k)}$ is the input feature of $a$ at the $k$-th layer, and is updated using con function to obtain the $k$-th-layer output as $\bm{a}^{(k+1)}=\textsc{con}(\bm{a}^{(k)}_{\mathcal{N}_{a}},\bm{a}^{(k)})$, similar for $o$ and $c$ with outputs $\bm{o}^{(k+1)}$ and $\bm{c}^{(k+1)}$. The input feature of $a,o,c$ at the first layer of GNN is the initialized node feature, e.g., $\bm{c}^{(0)}=\bm{h}_{c}^{t}$. The output features of $a,o$ after $K$ GNN layers $\bm{a}^{(K)},\bm{o}^{(K)}$, which have already fused their neighboring compositions’ features, and $c$’s output feature $\bm{c}^{(K)}$, which has aggregated the updated features of $a$ and $o$, are the final output of one L-Adapter, and will be inputted into the next computation unit of CLIP for the latter computation.
The design of V-Adapter

### v2

# Graph-guided Cross-composition Feature Disentanglement for Compositional Zero-shot Learning

@@ Abstract
Disentanglement of visual features of primitives (i.e., attributes and objects) has shown exceptional results in Compositional Zero-shot Learning (CZSL). However, due to the feature divergence of an attribute (resp. object) when combined with different objects (resp. attributes), it is challenging to learn disentangled primitive features that are general across different compositions. To this end, we propose the solution of cross-composition feature disentanglement, which takes multiple primitive-sharing compositions as inputs and constrains the disentangled primitive features to be general across these compositions. More specifically, we leverage a compositional graph to define the overall primitive-sharing relationships between compositions, and build a task-specific architecture upon the recently successful large pre-trained vision-language model (VLM) CLIP, with dual cross-composition disentangling adapters (called L-Adapter and V-Adapter) inserted into CLIP’s frozen text and image encoders, respectively. Evaluation on three popular CZSL benchmarks shows that our proposed solution significantly improves the performance of CZSL, and its components have been verified by solid ablation studies. Our code and data are available at: https://github.com/zhurunkai/DCDA.
Graph-guided Cross-composition Feature Disentanglement for Compositional Zero-shot Learning
Yuxia Geng1,2, Runkai Zhu2, Jiaoyan Chen3, Jintai Chen4, Xiang Chen5thanks: Corresponding Author with xiang_chen@nuaa.edu.cn, Zhuo Chen6, Shuofei Qiao6, Yuxiang Wang2, Xiaoliang Xu2, Sheng-Jun Huang5 1PowerChina Huadong Engineering Corporation Limited 2Hangzhou Dianzi University 3The University of Manchester 4The Hong Kong University of Science and Technology (Guangzhou) 5Nanjing University of Aeronautics and Astronautics 6Zhejiang University
1 Introduction
Compositional Zero-shot Learning (CZSL) aims to recognize novel attribute-object compositions by disentangling visual primitives from seen combinations, a capability crucial for scaling visual recognition systems Misra et al. (2017). For instance, a model trained on red tomato and green apple should infer green tomato through primitive recombination, despite never encountering this specific composition. This paradigm not only enables zero-shot generalization to exponentially many combinations Chen et al. (2023), but also advances vision-language understanding by requiring precise feature decomposition aligned with textual semantics Chen et al. (2025).
Early CZSL approaches establish shared embedding spaces to compare image features with composition embeddings Wei et al. (2019); Naeem et al. (2021); Mancini et al. (2022). Recent advances leverage CLIP’s visual-semantic alignment from large-scale pretraining Radford et al. (2021); Nayak et al. (2023), yet face inherent challenges: Attribute-object primitives exhibit strong visual entanglement—consider how red permeates all pixels of a red tomato. This entanglement hinders both primitive alignment and novel composition generalization. Current CLIP-based solutions address this through either disentangled text prompts Lu et al. (2023); Wang et al. (2023a) or vision adapters Zheng et al. (2024); Huang et al. (2024); Li et al. (2024), but crucially overlook the diversity of primitive manifestations across compositions. As Figure 1 illustrates, the visual realization of red varies significantly when combined with different objects (e.g., tomato vs. wine), exhibiting divergent color tones and spatial distributions.
Figure 1: Examples of divergent visual features of a primitive (e.g., an attribute red or an object tomato) across different compositions.
@@ 1 Introduction
Figure 2: $t$-SNE visualizations of disentangled attribute representations of images in the test set of MIT-States, learned by CAILA Zheng et al. (2024) and our DCDA. Solid and hollow circles represent images of seen and unseen compositions, respectively. Best viewed in color.
To quantify this challenge, we conduct feature-space analysis on MIT-States Isola et al. (2015) using CAILA Zheng et al. (2024), the current SOTA method. As visualized in Figure 2 (left), disentangled attributes like broken (purple circles) exhibit two critical limitations: (1) Features from the same attribute scatter widely with overlapping clusters (e.g., broken intermixed with cooked and mossy), indicating poor intra-class consistency; (2) This dispersion directly impacts generalization—compositions containing unseen broken objects become indistinguishable due to the attribute’s non-discriminative embeddings.
To overcome these limitations, we propose cross-composition feature aggregation through graph-guided learning. Our key insight is that effective primitive disentanglement requires cross-composition feature aggregation. Drawing inspiration from compositional graphs Naeem et al. (2021), we construct a tripartite graph (Figure 3a) connecting attributes, objects, and their compositions. This graph enables our Dual Cross-composition Feature Decomposing Adapters (DCDA), which enhance CLIP’s text and image encoders through complementary strategies: (1) L-Adapter propagates textual features across compositionally related nodes via GNNs, consolidating attribute/object semantics from multiple contexts. (2) V-Adapter employs cross-attention between primitive-sharing images (e.g., red tomato and red wine) to extract invariant visual patterns, augmented by a novel sampling strategy that weightedly selects compositions according to the attribute/object co-occurrence degrees (derived from our graph). When integrating L-Adapters and V-Adapters into multiple layers of CLIP’s text and image encoders, we retain the original parameters of CLIP to avoid overfitting, but inject the task-specific knowledge. Our contributions can be summarized below:
  • DCDA is the first systematic approach for cross-composition feature disentanglement in CLIP-based CZSL, explicitly addressing primitive diversity.
  • Dual adapter architecture leveraging compositional graphs for text-side aggregation and vision-side contrastive attention, enabling discriminative yet generalizable primitive representations.
  • DCDA achieves great performance on MIT-States and UT-Zappos (closed/open world), with 5.1%/7.3% gains over CAILA method [^1] Visualizations confirm tighter clustering of disentangled attributes (Figure 2, right), with extended object analyses in Appendix C..
2 Related Work
@@
Conventional CZSL methods are roughly divided into two groups. One is classifier-based which first trains two separate classifiers to predict an input image’s attribute and object labels, respectively, and then combines them to predict the compositional labels Misra et al. (2017); Nagarajan and Grauman (2018). The subsequent works further enhance the dependence of the attribute and object in a composition Li et al. (2020, 2022); Wang et al. (2023b). The other group is embedding-based which jointly represents attributes and objects to capture the dependence, and then aligns them with the images in a shared embedding space Wei et al. (2019); Karthik et al. (2022); Geng et al. (2021). In particular, Naeem et al. (2021) learn the joint representation through graph convolutional networks. There are also some works concerning the disentanglement of attribute and object features in the visual space Saini et al. (2022); Hao et al. (2023); Kim et al. (2023); Chen et al. (2022a, 2024b) or the label space Geng et al. (2022); Chen et al. (2022b, 2024a). However, all these methods have to learn the alignment between image features and text embeddings from scratch and are prone to overfit to the seen compositions. It is expected to derive pre-trained alignment knowledge from VLMs.
@@
CLIP-based CZSL. After pre-training using 400 million image-text pairs, CLIP can be applied to any visual classification task without fine-tuning by setting prompts like “a photo of [class]”, where “[class]” is filled with the name of the class to be recognized. Nayak et al. (2023) then had the first attempt to design prompt “a photo of [attribute] [object]” for CZSL, where “[attribute] [object]” are tunable tokens to teach CLIP how to compose attributes and objects.
@@
To stress the roles of individual primitives, Wang et al. (2023a) additionally set an attribute and an object prompt with only “[attribute]” or “[object]” tunable; Lu et al. (2023) make the whole prompt trainable and fuses the decomposed text features with the encoded (entangled) image features through a cross-modal fusion module. Different from these works focusing on optimizing the prompts, Zheng et al. (2024) propose to insert trainable adapters inside the frozen transformer layers to decompose and recompose the attribute and object features. With disentangled primitive features, Huang et al. (2024) establish three prediction branches and pulls a static class prompt to its dynamic images via a cross-modal traction module. Li et al. (2024) investigate the relative specificity of attributes when paired with different objects. In contrast to these methods, our method DCDA is more generalizable, with cross-composition knowledge injected and dual adapters inserted in CLIP’s image and text encoders.
@@
3 Methodology
CZSL Task Formulation. Let $\mathcal{D}_{tr}=\{(x,c)|x\in\mathcal{X}_{s},c\in\mathcal{C}_{s}\}$ be the training set, where $\mathcal{X}_{s}$ contains the training images and $\mathcal{C}_{s}$ is a set of seen compositional labels that are available during training. Each label is a tuple $c=(a,o)$ of an attribute class $a\in\mathcal{A}$ and an object class $o\in\mathcal{O}$. After training, the CZSL model can predict images of a set of new compositions $\mathcal{C}_{u}$ that are unseen during training, with $\mathcal{C}_{u}\cap\mathcal{C}_{s}=\emptyset$. Following previous works, we study generalized CZSL Purushwalkam et al. (2019), where images of seen and unseen compositions are both tested and the candidate label space includes both seen and unseen labels. The test set is thus denoted as $\mathcal{D}_{te}=\{(x,c)|x\in\mathcal{X}_{te},c\in\mathcal{C}_{te}\}$, where $\mathcal{X}_{te}=\mathcal{X}_{u}\cup\mathcal{X}^{\prime}_{s}$ with $\mathcal{X}^{\prime}_{s}\cap\mathcal{X}_{s}=\emptyset$, and $\mathcal{C}_{te}=\mathcal{C}_{u}\cup\mathcal{C}_{s}$. Notably, $\mathcal{C}_{s}$ and $\mathcal{C}_{u}$ share the same attribute set $\mathcal{A}$ and object set $\mathcal{O}$, CZSL assumes that each $a$ and $o$ has been trained before testing and only the composition $(a,o)\in\mathcal{C}_{u}$ is novel.
@@ 3.1 The design of L-Adapter
We first define the compositional graph. It consists of $N=|\mathcal{A}|+|\mathcal{O}|+|\mathcal{C}^{\prime}|$ nodes, including all the attributes, all the objects, and the compositions in the current computation (i.e., $\mathcal{C}^{\prime}=\mathcal{C}_{s}$ for training and $\mathcal{C}^{\prime}=\mathcal{C}_{te}$ for testing). Given these nodes, for each $c=(a,o)$, we connect $(a,o)$, $(a,c)$ and $(o,c)$ to form a triangle in the graph, as shown in Figure 3(a). For simplicity and efficiency, we keep all graph edges unweighted and undirected as in Naeem et al. (2021) and obtain a symmetric adjacency matrix $A\in\mathbb{R}^{N\times N}$ to store the graph structure, $A_{ij}=1$ if there is a connection between node $i$ and $j$ otherwise $A_{ij}=0$.
@@
Next, we exploit multiple GNN layers to propagate features among graph nodes following the graph structure defined in $A$. Formally, as Figure 3(a)’s top shows, for each $c=(a,o)$, three identical AGG functions are parallelly applied to aggregate neighborhood features for nodes $a,o,c$ at the $k$-th GNN layer, $k\in\{0,...,K-1\}$, as:
$\bm{a}^{(k)}_{\mathcal{N}_{a}}=\text{AGG}^{(k)}(\{\bm{c}^{(k)}_{i}|c_{i}\in\mathcal{N}_{a}^{c}\},\{\bm{o}^{(k)}_{i}|o_{i}\in\mathcal{N}_{a}^{o}\})$ (1)
$\bm{o}^{(k)}_{\mathcal{N}_{o}}=\text{AGG}^{(k)}(\{\bm{c}^{(k)}_{j}|c_{j}\in\mathcal{N}_{o}^{c}\},\{\bm{a}^{(k)}_{j}|a_{j}\in\mathcal{N}_{o}^{a}\})$ (2)
$\bm{c}^{(k)}_{c=(a,o)}=\text{AGG}^{(k)}(\bm{a}^{(k)},\bm{o}^{(k)})$ (3)
where $\mathcal{N}_{a}^{c}$ (resp. $\mathcal{N}_{o}^{c}$) denotes the composition neighbor set of $a$ (resp. $o$) on the graph, and $\mathcal{N}_{a}^{o}$ (resp. $\mathcal{N}_{o}^{a}$) includes the objects (resp. attributes) that compose the compositions in $\mathcal{N}_{a}^{c}$ (resp. $\mathcal{N}_{o}^{c}$) together with $a$ (resp. $o$). Considering the example in Figure 3(a) where $c=(a,o)$ is red tomato, $\mathcal{N}_{a}^{c}$ includes compositions like red apple and $\mathcal{N}_{a}^{o}$ includes objects like apple. While the neighbor set of each $c$ only contains its primitives $a$ and $o$. $\bm{a}^{(k)}$ is the input feature of $a$ at the $k$-th layer, and is updated using CON function to obtain the $k$-th-layer output as $\bm{a}^{(k+1)}=\text{CON}(\bm{a}^{(k)}_{\mathcal{N}_{a}},\bm{a}^{(k)})$, similar for $o$ and $c$ with outputs $\bm{o}^{(k+1)}$ and $\bm{c}^{(k+1)}$. The input feature of $a,o,c$ at the first layer of GNN is the initialized node feature, e.g., $\bm{c}^{(0)}=\bm{h}_{c}^{t}$. The output features of $a,o$ after $K$ GNN layers $\bm{a}^{(K)},\bm{o}^{(K)}$, which have already fused their neighboring compositions’ features, and $c$’s output feature $\bm{c}^{(K)}$, which has aggregated the updated features of $a$ and $o$, are the final output of one L-Adapter, and will be inputted into the next computation unit of CLIP for the latter computation.
3.2 The design of V-Adapter